OpenStudy (anonymous):

How do you write an equation of a line with the x intercept of 5 and a y intercept of 6? Can be left in point intercept form for ease! :)

Directrix (directrix):

An x-intercept has coordinates (5,0) because to "get" on the x-axis, your y-coordinate has to be 0. By similar reasoning, the y-intercept has coordinates (0,6). The slope of the line is (0 - 6) / (5 - 0) = -6/5. Use either point for (x1, y1). Using the y-intercept: y - 6 = (-6/5) ( x - 0) y - 6 = (-6/5) x + 0 y = (-6/5) x + 6